how is calculating 25% of $15 the same as calcuating 15% of $25? how are the calculaton different?sketch number line to illustrate ur answer?
1
Expert's answer
2012-11-09T03:32:59-0500
25% of $15 would be 15 * (25/100) = 15 * 0,25 = 3,75, while& 15% of $25 = 25 * (15/100) = 25 * 0,15 = 3,75